1. Calculate the cash payback period for each of the products listed below. The investment required is $400,000 for both. Be sure to show all work to receive full credit and label each answer clearly.

Year              Product A                     Product B
1                      70,000                         70,000
2                      80,000                         70,000
3                      90,000                         70,000
4                     100,000                         70,000
5                      60,000                         70,000
6                      50,000                         70,000
7                      40,000                         70,000

Total               490,000                        490,000

2. Production and sales estimates for May for the Finneaty Co. are as follows:

Estimated inventory (units), March 1

17,500

Desired inventory (unit), March 31

19,300

Expected sales volume (units):

  Area W

4,200

  Area X

7,000

  Area Y

9,000

Unit sales price

$15

The number of units expected to be sold in May is:

22,000

1,800

23,800

20,200

3. Ecology Company sells a biodegradable product called Run-All and has the following sales on account for the first four months of the current year:

                                                Jan                   Feb                  March              April

Sales in units:                            2,000               2,000               3,000               4,000

The company is preparing a cash budget for April only. Based on past history accounts are collected 50% in the current month, 20% from the prior month, and 20% from 2 months back. Ten percent is considered uncollectible. How much cash will be collected in April?

8700

4000

3000

2000

4. Explain the difference between fixed and variable costs per unit AND in total. Provide an example to support your explanation
Answer #1

Answer 1:

Payback period refers to the amount of time it takes to recover the cost of an investment. In the given question amount of Investment is $4,00,000, calculation of payback period will be as follows:

Where the cash inflow is uneven the calculation of payback period will be as follows-

Payback period = A+(B/C)

Where,
A is the last period number with a negative cumulative cash flow;
B is the absolute value (i.e. value without negative sign) of cumulative net cash flow at the end of the period A; and
C is the total cash inflow during the period following period A

Year Cash inflow of product A Cumulative Cash inflow of product A

0 -400000 -400000

1 70000 -330000

2 80000 -250000

3 90000 -160000

4 100000 -60000

5 60000 0

6 50000 50000

7 40 90000

Payback period = 4+(60000/60000)

=4 years

Where the cash inflow is even the calculation of payback period will be as follows-

Payback period = Cash outflow/Cash inflow

For product B the cash inflows are even during the years, so payback period will be as follows-

Payback period = 400000/70000

= 5.71 years

Answer 3:

As per past history collection of accounts are as follows:

Current month = 50%

Prior month = 20%

2 months back = 20%

Collection in April month is as follows:

50% of April's sales = 4000 * 50% = 2000

20% of March's sales = 3000*20% = 600

20% of February's sales = 2000*20% = 400

Total collection in April = 2000+600+400 = 3000

Answer 4:

Fixed cost is the cost which remains same regardless the volume produced, even it will be incur whether there is any production or not. For example depreciation of assets, insurance etc.

Variable cost is the cost which changes with the change in output, it is incur only when the units produced. For example Material and Labour cost, packing cost etc.